# A Scalable Inter-edge Correlation Modeling in CopulaGNN for Link Sign Prediction

### v3

@@ 1 Introduction
A signed graph refers to a graph where edges can be either positive (++) or negative (-). In the real world, a variety of systems containing positive and negative relationship between entities can be modeled as a signed graph; e.g., social networks (friendly/hostile) or content recommendation (like/dislike) (Liao et al., 2018; Chen et al., 2024). On a signed graph, we are interested in the link sign prediction task, aiming at predicting the sign of unobserved edges, given the partially observed positive and negative edges in a graph (Agrawal et al., 2013).
@@
However, these additional auxiliary structures often lead to slow convergence, and in some cases, significantly inefficient use of the memory space. In contrast to adding auxiliary operations, we shift our attention to a more fundamental aspect: statistical dependency among edges. Instead of assuming that the adjacent nodes are similar, we assume that adjacent edges connected via a common node are not informatively independent; that is, they are either in accordance with or in the opposite to each other. This relaxation naturally extends the idea of GNNs from an unsigned graph to a signed one. Note that our assumption is edge-focused since we target the link sign prediction problem.
@@ 3.1 Joint Edge Label Distribution as Link Sign Prediction Model
It is noteworthy that while the correlation matrix $\mathbf{R}\in[-1,1]^{n\times n}$ is usually huge, its structure can be represented as the Gramian form thanks to the properties of the normalization function $\nu(\cdot)$. For this reason, the Gaussian copula $\mathcal{C}$ can be implemented with a low-rank multivariate Gaussian distribution without imposing a memory bottleneck. Further details are provided in Appendix C.
@@ 3.2 Model Training
Generalization to Unobserved Edges. Although we train our model using only the observed edges in the graph, the SNEA encoder in Fig. 1 generalizes this learned knowledge to all nodes. Crucially, we adhere to the standard setting of link sign prediction. Under this standard setting, the objective is to determine the sign of a relationship given the premise that the edge structurally exists. Therefore, an unobserved edge under our context refers to an existing edge with its sign unknown, distinct from the regular link prediction task. This formulation allows us to directly compute not only their edge embeddings but also their marginal distributions via the learned linear projections. Consequently, our model $\mathcal{H}$ is applicable to the entire $\mathbf{R}$ as in Eq. (9), not just to its observed portion $\mathbf{R}_{00}$.
@@ 3.3 Efficient Inference
To predict the sign of an unseen edge, we may perform inference by directly sampling from the fitted conditional distribution, given the known labels in $\mathcal{E}_{\mathrm{obs}}$. Relying on a fundamental property that a conditional distribution of a Gaussian given another Gaussian is also a Gaussian, we may perform a naive sampling analytically from our Gaussian copula. Specifically, for $\mathbf{z}_{\mathrm{miss}}:=(\Phi^{-1}(u_{m+1}),\Phi^{-1}(u_{m+2}),\cdots,\Phi^{-1}(u_{n}))$, the conditional distribution from which we sample is given by:
@@ 5.1 Experimental Setup
We compare our method with state-of-the-art baselines, including GCN (Kipf and Welling, 2017), SGCN (Derr et al., 2018), SNEA (Li et al., 2020), SDGNN (Huang et al., 2021), TrustSGCN (Kim et al., 2023), SLGNN (Li et al., 2023), SE-SGformer (Li et al., 2025), and SGAAE (Nakis et al., 2025), on multiple real-world datasets including BitcoinAlpha (Kumar et al., 2016), BitcoinOTC (Kumar et al., 2018), WikiElec (Leskovec et al., 2010a), WikiRfa (West et al., 2014), SlashDot, and Epinions (Leskovec et al., 2010b). We report AUC and (Macro) F1 as our primary metrics, consistent with previous studies. All results are averaged over 10 distinct data splits to ensure statistical reliability. Appendix F provides further experimental details and supplementary results.
@@ 6 Related Work
Signed Graph Neural Networks. A key challenge in signed graph learning is that negative edges violate the graph homophily assumption. Early work, such as SGCN (Derr et al., 2018) and SDGNN (Huang et al., 2021), addressed this by incorporating balance theory (Heider, 1946) or status theory (Easley and Kleinberg, 2010). Subsequently, SiGAT (Huang et al., 2019) and SNEA (Li et al., 2020) further enhanced representation quality by applying the graph attention (Velickovic et al., 2018). MSGNN (He et al., 2022) and SLGNN (Li et al., 2023) refines representation by emphasizing low-frequency signals through the utilization of graph spectral filtering (Defferrard et al., 2016; Kim et al., 2025). More recent efforts, such as TrustSGCN (Kim et al., 2023), SE-SGformer (Li et al., 2025), and SGAAE (Nakis et al., 2025), have shifted focus toward reliability and explainability, using trust-aware propagation and explainable architectures. Despite these advances, a common limitation across these methods is their separate handling of positive and negative edges during aggregation, which overlooks the statistical dependencies that may exist between them.
@@ Acknowledgments
This research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2025-25399604).
@@ Appendix A Relaxed Bernoulli Distribution
$\frac{dF}{dx}={{\frac{atx^{t-1}(1-x)^{t-1}}{(a(1-x)^{t}+x^{t})^{2}}}}=f(x;a,t).$

### v4

@@ 1 Introduction
A signed graph refers to a graph where edges can be either positive (++) or negative (-). In the real world, a variety of systems containing positive and negative relationship between entities can be modeled as a signed graph; e.g., social networks (friendly/hostile) or content recommendation (like/dislike) (Liao et al., 2018; Kim et al., 2024; Chen et al., 2024; Kim et al., 2025b). On a signed graph, we are interested in the link sign prediction task, aiming at predicting the sign of unobserved edges, given the partially observed positive and negative edges in a graph (Agrawal et al., 2013).
@@
However, these additional auxiliary structures often lead to slow convergence, and in some cases, significantly inefficient use of the memory space. In contrast to adding auxiliary operations, we shift our attention to a more fundamental aspect: statistical dependency among edges. Instead of assuming that the adjacent nodes are similar, we assume that adjacent edges connected via a common node are not informationally independent; that is, they are either in accordance with or in the opposite to each other. This relaxation naturally extends the idea of GNNs from an unsigned graph to a signed one. Note that our assumption is edge-focused since we target the link sign prediction problem.
@@ 3.1 Joint Edge Label Distribution as Link Sign Prediction Model
It is noteworthy that while the correlation matrix $\mathbf{R}\in[-1,1]^{n\times n}$ is usually huge, its structure can be represented as the Gramian form thanks to the properties of the normalization function $\nu(\cdot)$. For this reason, the Gaussian copula $\mathcal{C}$ can be implemented with a low-rank multivariate Gaussian distribution without imposing a memory bottleneck. (see Appendix C).
@@ 3.2 Model Training
Generalization to Unobserved Edges. Although we train our model using only the observed edges in the graph, the signed graph encoder (implemented as SNEA (Li et al., 2020) in Fig. 1) generalizes this learned knowledge to all nodes. Crucially, we adhere to the standard setting of link sign prediction. Under this standard setting, the objective is to determine the sign of a relationship given the premise that the edge structurally exists. Therefore, an unobserved edge under our context refers to an existing edge with its sign unknown, distinct from the regular link prediction task. This formulation allows us to directly compute not only their edge embeddings but also their marginal distributions via the learned linear projections. Consequently, our model $\mathcal{H}$ is applicable to the entire $\mathbf{R}$ as in Eq. (9), not just to its observed portion $\mathbf{R}_{00}$.
@@ 3.3 Efficient Inference
To predict the sign of an unseen edge, we may perform inference by directly sampling from the fitted conditional distribution, given the known labels in $\mathcal{E}_{\mathrm{obs}}$. Relying on a fundamental property that a conditional distribution of a Gaussian given another Gaussian is also a Gaussian, we may perform a naive sampling analytically from our Gaussian copula. Specifically, for $\mathbf{z}_{\mathrm{miss}}:=(\Phi^{-1}(u_{m+1}),\Phi^{-1}(u_{m+2}),\cdots,\Phi^{-1}(u_{n}))$, the conditional Gaussian distribution from which we sample is given by:
@@ 5.1 Experimental Setup
We compare our method with state-of-the-art baselines, including GCN (Kipf and Welling, 2017), SGCN (Derr et al., 2018), SNEA (Li et al., 2020), SDGNN (Huang et al., 2021), TrustSGCN (Kim et al., 2023), SLGNN (Li et al., 2023), SE-SGformer (Li et al., 2025), and SGAAE (Nakis et al., 2025), on multiple real-world datasets including BitcoinAlpha (Kumar et al., 2016), BitcoinOTC (Kumar et al., 2018), WikiElec (Leskovec et al., 2010a), WikiRfa (West et al., 2014), SlashDot, and Epinions (Leskovec et al., 2010b). We report AUC and (Macro) F1 as our primary metrics, consistent with previous studies. All results are averaged over 10 distinct data splits to ensure statistical reliability. Further experimental details and supplementary results are provides in Appendix F.
@@ 6 Related Work
Signed Graph Neural Networks. A key challenge in signed graph learning is that negative edges violate the graph homophily assumption. Early work, such as SGCN (Derr et al., 2018) and SDGNN (Huang et al., 2021), addressed this by incorporating balance theory (Heider, 1946) or status theory (Easley and Kleinberg, 2010). Subsequently, SiGAT (Huang et al., 2019) and SNEA (Li et al., 2020) further enhanced representation quality by applying the graph attention (Velickovic et al., 2018). MSGNN (He et al., 2022) and SLGNN (Li et al., 2023) refines representation by emphasizing low-frequency signals through the utilization of graph spectral filtering (Defferrard et al., 2016; Kim et al., 2025a). More recent efforts, such as TrustSGCN (Kim et al., 2023), SE-SGformer (Li et al., 2025), and SGAAE (Nakis et al., 2025), have shifted focus toward reliability and explainability, using trust-aware propagation and explainable architectures. Despite these advances, a common limitation across these methods is their separate handling of positive and negative edges during aggregation, which overlooks the statistical dependencies that may exist between them.
@@ Acknowledgments
This work was supported by SOFT Foundry Institute at SNU, Youlchon Foundation, National Research Foundation of Korea (NRF) grants (RS-2021-NR05515, RS-2024-00336576, RS-2023-0022663, RS-2025-25399604) and Institute for Information & communication Technology Planning & evaluation (IITP) grants (RS-2022-II220264, RS-2024-00353131) by the government of Korea.
@@ Appendix A Relaxed Bernoulli Distribution
$\frac{dF}{dx}={{\frac{atx^{t-1}(1-x)^{t-1}}{(a(1-x)^{t}+x^{t})^{2}}}}=f_{\mathrm{simple}}(x;a,t).$